Write answer without exponents
(1/32)^2/5
64^-2/3

Answer by RAY100(1637)

(1/32)^2/5 = {(1/32)^1/5}^2 =1/2^2 = 1/4
.
.
(64)^-2/3 = 1/(64)^2/3 - 1/ {(64)^1/3}^2 = 1/(4)^2 = 1/16
.